OpenStudy (anonymous):

What is the value of (8 – 3)!? A:336 B:5 C:120 D:40,314 WILL GIVE MEDAL!!!PLZ HELPPPP

OpenStudy (raden):

8 – 3 = 5 so, (8 – 3)! = 5! = 5 * 4 * 3 * 2 * 1 = ...

OpenStudy (anonymous):

120?

OpenStudy (raden):

yup, correct